After the release of radioactive material into the atmosphere from a nuclear power plant in a country in 1987​, the hay in that country was contaminated by a radioactive isotope​ (half-life 6 ​days). If it is safe to feed the hay to cows when 88​% of the radioactive isotope is​ gone, how long did the farmers need to wait to use this​ hay?
Answer by Boreal(15235)

P=Poe^(-kt)
1/2=e^(-6k)
ln of both sides
-0.693=-6k
k=0.1155
0.12=e^(-.1155*t)
ln of both sides
-2.12=-.1155t
t=18.35 days.
Check.
half gone in 6 days
three-quarters gone in 12 days
seven-eights gone i 18 days. (0.125)